OpenStudy (anonymous):

factor r^2-r-6 is the answer r=2 r=3?

OpenStudy (cwrw238):

no when multiplied the 2 numbers must = -6

OpenStudy (anonymous):

so then they should be negatives?

OpenStudy (cwrw238):

one should be negative and the other positive and their sum must be -1 ( because middle term is -r)

OpenStudy (anonymous):

so negative 2 and positive 3?

OpenStudy (cwrw238):

no that would give -2 + 3 = +1

OpenStudy (anonymous):

oh ok then, negative 3 and positive 2 sorry

OpenStudy (cwrw238):

yes r^2 - r - 6 = r^2 -3r + 2r - 6 = r(r - 3) + 2(r - 3) = (r - 3)(r + 2) - these are the factors not r = 3 and r = -2 which are the zeros of the function

OpenStudy (anonymous):

ok, thanks